L^{A}TEX News

Issue 4, December 1995

Welcome to L^{A}TEX News 4

An issue of L^{A}TEX News will accompany every future release of L^{A}TEX. It will tell you about important events, such as major bug fixes, newly available packages, or any other L^{A}TEX news. This issue accompanies the fourth release of L^{A}TEX 2ε.

L^{A}TEX getting smaller

The last release in June started a trend of L^{A}TEX becoming smaller, we are pleased to announce that this has continued with this release. In particular the experimental ‘autoload’ version described in autoload.txt is much smaller as more parts of L^{A}TEX are autoloaded.

New ‘concurrent’ docstrip

The time taken to ‘unpack’ this release from the documented sources should be much reduced (roughly half the time, depending on installation conditions). This is due to an improved version of the docstrip program that has been contributed by Marcin Woliński. This can write up to 16 files at once. The previous version could only write one file at a time which meant that it was very slow when producing many small files from the same source file as the source needed to be re-read for each file written.

New T1 encoded fonts

This year Jörg Knappen has completed a new release of the ‘Cork’ (T1) encoded Computer Modern fonts: the dc fonts release 1.2.

This release of the dc fonts fixes many bugs (including the missing ?` (¿) and !` (¡) ligatures) and improves the fonts in many other ways. It is strongly recommended that you upgrade as soon as possible if currently you are using the old dc fonts, release 1.1 or earlier. The new fonts are available from the CTAN

archives, in tex-archive/fonts/dc.

The names of the font files are different. This does not affect L^{A}TEX documents but does affect the installation procedure as it assumes that you have the new fonts, and will write suitable ‘fd’ files for those fonts. If you have not yet upgraded your dc fonts then, after unpacking the distribution, you must latex olddc.ins to produce ‘fd’ files for the old dc fonts. This must be done before the format is made. Running the test document at ltxcheck.tex the end of the installation will inform you if the wrong set of ‘fd’ files has been installed.

Note that this change does not affect the standard ‘OT1’ Computer Modern fonts that L^{A}TEX uses by default.

More robust commands

The commands \cite and \sqrt are now robust.

Although most commands with optional arguments are fragile, as documented, such commands defined using the second optional argument of \newcommand and its derivatives are now robust.

New Interface to building ‘extension’ classes

The mechanism provided by \DeclareOption,

                                                   \ProcessOptions and \LoadClass has proved to be a

powerful and expressive means of defining one class in terms of another ‘base’ class. However there have been some requests to simplify the declaration of the common case where you want the ‘base’ class to be called with all the options that were specified to the extension class. This is now provided by the new

                                                   command \LoadClassWithOptions. A similar
                                                   command \RequirePackageWithOptions is provided

for package use. More details of this feature are

                                                   provided in clsguide.tex and ltclass.dtx.

More Input Encodings

The experimental inputenc package allows a more natural style of input of accented and other characters. Three new input encodings are now supported.

• ansinew the Windows ansi encoding, as used in Microsoft Windows 3.x.

• cp437de a variant of cp437, which uses ß rather than β in the appropriate slot.

• next the encoding used on Next computers.

Further information

For more information on TEX and L^{A}TEX, get in touch with your local TEX Users Group, or the international TEX Users Group, 1850 Union Street, #1637, San Francisco, CA 94123, USA, Fax: +1 415 982 8559, EMail: tug@tug.org. The L^{A}TEX home page is

                                                   http://www.tex.ac.uk/ctan/latex/ and contains

links to other WWW resources for L^{A}TEX.

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 1995, all rights reserved.